# Boosting Robustness by Clipping Gradients in Distributed Learning

### v1

@@ Abstract
Robust distributed learning consists in achieving good learning performance despite the presence of misbehaving workers. State-of-the-art (SOTA) robust distributed gradient descent (Robust-DGD) methods, relying on robust aggregation, have been proven to be optimal: Their learning error matches the lower bound established under the standard heterogeneity model of $(G,B)$-gradient dissimilarity. The learning guarantee of SOTA Robust-DGD cannot be further improved when model initialization is done arbitrarily. However, we show that it is possible to circumvent the lower bound, and improve the learning performance, when the workers’ gradients at model initialization are assumed to be bounded. We prove this by proposing pre-aggregation clipping of workers’ gradients, using a novel scheme called adaptive robust clipping (ARC). Incorporating ARC in Robust-DGD provably improves the learning, under the aforementioned assumption on model initialization. The factor of improvement is prominent when the tolerable fraction of misbehaving workers approaches the breakdown point. ARC induces this improvement by constricting the search space, while preserving the robustness property of the original aggregation scheme at the same time. We validate this theoretical finding through exhaustive experiments on benchmark image classification tasks.
@@ 1 Introduction
Distributed machine learning, a.k.a. federated learning, has emerged as a dominant paradigm to cope with the increasing computational cost of learning tasks, mainly due to growing model sizes and datasets [25]. Worker machines, holding each a fraction of the training dataset, collaborate over a network to learn an optimal common model over the collection of their datasets. Workers typically collaborate with the help of a central coordinator, that we call server [36]. Besides scalability, distributed learning is also helpful in preserving data privacy, since the workers do not have to share their local datasets during the learning.
@@
Conventional distributed learning algorithms are known to be vulnerable to misbehaving workers that could behave unpredictably [7, 25, 18]. Misbehavior may result from software and hardware bugs, data poisoning, or malicious players controlling part of the network. In the parlance of distributed computing, misbehaving workers are referred to as Byzantine [32]. Due to the growing influence of distributed learning in critical public-domain applications, such as healthcare [38] and finance [34], the problem of robustness to misbehaving workers, a.k.a. robust distributed learning, has received significant attention in the past [7, 44, 12, 14, 27, 17, 3, 13].
@@
Robust distributed learning algorithms primarily rely on robust aggregation, such as coordinate-wise trimmed mean (CWTM) [44], geometric median (GM) [10] and multi-Krum (MK) [7]. Specifically, in a robust distributed gradient descent (Robust-DGD) method, the server aggregates the workers’ local gradients using a robust aggregation rule, instead of simply averaging them. This protects the learning procedure from erroneous gradients sent by misbehaving workers. Recent work has made significant improvements over these aggregation rules: by incorporating a pre-aggregation step such as bucketing [27, 17] and nearest-neighbor mixing (NNM) [3] to tackle gradient dissimilarity that results from data heterogeneity. The learning guarantee of a resulting Robust-DGD method has been proven to be optimal [4], i.e., it cannot be improved unless additional assumptions are made, under the standard heterogeneity model of $(G,B)$-gradient dissimilarity [28].
Despite the theoretical tightness of state-of-the-art (SOTA) Robust-DGD, we demonstrate that its performance can be improved in pragmatic cases wherein model initialization incurs bounded errors on workers’ local data. We do so by proposing clipping of workers’ gradients prior to their aggregation (at the server) using a novel clipping scheme, named adaptive robust clipping (ARC). We show that incorporating ARC provably improves the robustness of Robust-DGD when the local gradients at the initial model estimate are bounded. The factor of improvement is prominent when the tolerable fraction of misbehaving workers approaches the breakdown point, [^1] The breakdown point is the minimum fraction of misbehaving workers that can break the system, i.e., makes it impossible to guarantee any bound on the learning error [18]. established in [4] under $(G,B)$-gradient dissimilarity. In general, when the local gradients can be arbitrarily large at initialization, ARC recovers the original learning guarantee of Robust-DGD.
Main results & contributions.
We consider a system comprising $n$ workers and a server. The goal is to tolerate up to $f$ misbehaving (a.k.a. adversarial) workers. The impact of data heterogeneity is characterized by the standard definition of $(G,B)$-gradient dissimilarity [28], recalled in Section 2.
Figure 1: Worst-case maximal accuracies across several types of misbehaviour (see Section 6) of CWTM $\circ$ NNM and GM $\circ$ NNM, with ARC and without (i.e., no clipping), for a heterogeneously distributed MNIST dataset with $10$ honest workers. Heterogeneity is simulated using Dirichlet [23] distribution of parameter $\alpha=0.5$.
@@
(1) Adaptive robust clipping (ARC). We propose ARC, wherein prior to aggregating the gradients, the server clips the largest $k\coloneqq\left\lfloor{2(f/n)(n-f)}\right\rfloor$ gradients using a clipping parameter given by the (Euclidean) norm of the $(k+1)$-th largest gradient. We prove 2 crucial properties: ARC (i) preserves the robustness guarantee of original aggregation and (ii) constrains an adversarial gradient’s norm by that of an honest (non-adversarial) worker’s gradient.
(2) Improved learning error. Under $(G,B)$-gradient dissimilarity, [4] proves a lower bound on the learning error in the presence of adversarial workers. We show, however, that this lower bound can be circumvented under the assumption that the honest workers’ gradients are bounded at model initialization. Specifically, under this assumption, we prove that incorporating ARC in Robust-DGD improves the learning error beyond the lower bound when $f/n$ approaches the breakdown point. Thanks to its two aforementioned properties, ARC induces this improvement by effectively reducing the heterogeneity from $(G,B)$ to $(G+B\rho,0)$-gradient dissimilarity, where $\rho$ depends linearly on the norm of workers’ gradients at model initialization. We remark that the above assumption on model initialization is often satisfiable in practice [15].
@@
(3) Empirical validation. We conduct experiments on MNIST [11], Fashion-MNIST [42], and CIFAR-10 [31], under several data heterogeneity and adversarial regimes. Our results show that ARC boosts the performance of SOTA robust aggregation methods, especially when considering a large number of adversarial workers. A preview of our results on MNIST is shown in Figure 1. We make similar observations on the effectiveness of ARC in scenarios of high heterogeneity (see Figure 3).
Related work.
Gradient clipping is a well-known technique for tackling exploding gradients in deep learning [16]. It has been extensively analyzed in the centralized setting [46, 45, 30], with applications to differential privacy [39]. There is also some work analyzing the impact of clipping in non-robust distributed learning [47, 29]. It has been shown that the clipping of local gradients causes distributed gradient descent to diverge. This motivates the need for an analysis of the clipping technique in our adversarial setting, and also hints at the inadequacy of clipping with a fixed threshold.
In the context of robust distributed learning, prior work has proposed iterative clipping for robust aggregation [26]. The analysis, however, cannot be applied to study the impact of pre-aggregation clipping that we consider. An adaptive gradient clipping scheme was also proposed for robust aggregation in [20, 33]. However, the robustness guarantees are only provided for strongly convex loss functions and under a specific redundancy condition that is inadequate for our non-convex machine learning setting. Other work [37, 14, 3] has only used static pre-aggregation clipping to enhance the empirical robustness of the proposed algorithms, but fail to provide any theoretical guarantees. Recent work [35] has proposed clipping of the gradient differences, in conjunction with variance reduction, to tackle partial worker participation when adversarial workers can form a majority. Our focus is however different, as we consider the impact of pre-aggregation clipping on the robustness obtained by a general class of aggregation rules.
Lastly, prior work [19, 48, 5] has also considered the challenge of privacy preservation alongside robustness in distributed learning. In this line of work, the gradients are assumed to have bounded norms in order to control the sensitivity of the algorithm, as is common in differential privacy (DP) [1]. This is often ensured in practice by clipping the gradients using a static clipping parameter. Our findings suggest that it is better to use adaptive clipping instead of static clipping. Nevertheless, thoroughly analysing ARC in the context of DP remains an interesting research direction.
Paper outline.
Section 2 presents the problem statement. Section 3 presents the background on robust distributed learning, while Section 4 introduces ARC and its useful properties. Section 5 presents the convergence analysis of Robust-DGD with ARC, and characterizes the improvement in its robustness. Section 6 presents our empirical results. Lastly, Section 7 presents concluding remarks. Appendix A presents the computational complexity of ARC. Appendices B and C contain proofs for results presented in Sections 4 and 5, respectively. Appendix D describes the experimental setup, while Appendices E and F include additional experimental results.
2 Problem Statement
We consider the problem of distributed learning in the server-based architecture. The system comprises $n$ workers represented by $w_{1},\dots,w_{n}$, that collaborate with the help of a trusted server. The workers hold local datasets $\mathcal{D}_{1},\dots,\mathcal{D}_{n}$ respectively, each composed of $m$ data points from an input space $\mathcal{Z}$. Specifically, for any $i\in[n]$, $\mathcal{D}_{i}\coloneqq\{z_{1}^{(i)},\dots,z_{m}^{(i)}\}\subset\mathcal{Z}^{m}$. For a given model parameterized by vector $\theta\in\mathbb{R}^{d}$, $d$ being the number of trainable parameters in the model, each worker $w_{i}$ incurs a loss given by the loss function $\mathcal{L}_{i}{(\theta)}\coloneqq\frac{1}{m}\sum_{k=1}^{m}\ell(\theta,z_{k}^{(i)})\enspace$, where $\ell:\mathbb{R}^{d}\times\mathcal{Z}\rightarrow\mathbb{R}$ is the point-wise loss. We make the following standard assumptions [8]: (i) the point-wise loss function $\ell$ is differentiable with respect to $\theta$. (ii) For all $i\in[n]$, $\mathcal{L}_{i}$ is $L$-Lipschitz smooth, i.e., there exists $L\in\mathbb{R}^{+}$ such that all $\theta,\theta^{\prime}\in\mathbb{R}^{d}$, $\left\lVert{\nabla\mathcal{L}_{i}(\theta{})-\nabla\mathcal{L}_{i}(\theta{}^{\prime})}\right\rVert\leq L\left\lVert{\theta{}-\theta{}^{\prime}}\right\rVert$. In the ideal setting where all workers are assumed to be honest, i.e., follow the prescribed algorithm correctly, the server aims to minimize the global average loss function given by $\mathcal{L}(\theta{})\coloneqq\frac{1}{n}\sum_{i=1}^{n}\mathcal{L}_{i}(\theta{})$. However, this learning objective is rendered vacuous when some workers could be adversarial, described in the following.
@@ 2 Problem Statement
Robust distributed learning. A robust distributed learning algorithm aims to output a good model despite the presence of some adversarial (a.k.a., Byzantine [32]) workers in the system [40, 44, 3]. Specifically, the goal is to design a distributed learning algorithm that can tolerate up to $f$ adversarial workers, of a priori unknown identity, out of $n$ workers. Adversarial workers need not follow a prescribed algorithm, and can send arbitrary information to the server. In the context of distributed gradient descent, adversarial workers can send incorrect gradients to the server [6, 2, 43, 27]. We let $\mathcal{H}\subseteq[n]$, with $\left\lvert{\mathcal{H}}\right\rvert=n-f$, denote the set of honest workers that do not deviate from an algorithm. The objective of the server is to minimize the honest workers’ average loss function given by
@@
While we can find a minimum of $\mathcal{L}_{\mathcal{H}}(\theta)$ when $\mathcal{L}_{\mathcal{H}}$ is convex, in general however the loss function is non-convex, and the optimization problem is NP-hard [9]. Therefore, we aim to find a stationary point of $\mathcal{L}_{\mathcal{H}}$ instead, i.e., $\theta^{*}$ such that $\left\lVert{\nabla\mathcal{L}_{\mathcal{H}}(\theta^{\ast})}\right\rVert=0$. Formally, we define robustness to adversarial workers by $(f,\epsilon)$-resilience [3].
@@
If a distributed learning algorithm $\mathcal{A}$ is $(f,\varepsilon)$-resilient then it can tolerate up to $f$ adversarial workers. Thus, $f/n$ is referred to as the maximum tolerable fraction of adversarial workers for $\mathcal{A}$. It is standard in robust distributed learning to design an algorithm using $f$ as a parameter [40, 7, 44, 27, 17, 3].
Data heterogeneity. In the context of distributed learning, data heterogeneity is characterized by the following standard notion of $(G,B)$-gradient dissimilarity [41, 28, 27, 17, 4].
@@
3 Relevant Background
We present in this section useful precursors on robust distributed learning.
General limitations on robustness. Note that $(f,\varepsilon)$-resilience is impossible (for any $\varepsilon$) when $f/n\geq 1/2$ [20]. Moreover, even for $f/n<1/2$, it is generally impossible to achieve $(f,\varepsilon)$-resilience for arbitrarily small $\varepsilon$ due to the disparity amongst the local datasets $\mathcal{D}_{i},\;i\in\mathcal{H}$ (a.k.a., data heterogeneity) [33]. Henceforth, we assume that $n>2f$ by default. Under $(G,B)$-gradient dissimilarity, we have the following lower bound on the achievable resilience.
Lemma 3.1 (Non-convex extension of Theorem 1 in [4]).
Under $(G,B)$-gradient dissimilarity, a distributed learning algorithm $\mathcal{A}$ satisfies $(f,\varepsilon)$-resilience only if $\frac{f}{n}<\frac{1}{2+B^{2}}$ and $\varepsilon\geq\frac{1}{4}\cdot\frac{f}{n-(2+B^{2})f}G^{2}$.
For a given distributed learning problem, the minimum fraction of adversarial workers that renders any distributed learning algorithm ineffective is referred the breakdown point [18, 4]. Thus, according to Lemma 3.1, there exists a distributed learning problem satisfying $(G,B)$-gradient dissimilarity whose breakdown point is given by $1/(2+B^{2})$.
In the ideal setting, i.e., when all the workers are honest, the server can execute the classic distributed gradient descent (DGD) method in collaboration with the workers to find a stationary point of the global average loss function. In DGD, the server maintains the model, which is updated periodically using the average of the gradients computed by the workers on their local datasets. The averaging of the gradients is rendered ineffective when some workers are adversarial, i.e., they can send malicious vectors posing as their gradients to the server [18]. To remedy this, averaging is replaced with robust aggregation, e.g., coordinate-wise trimmed mean (CWTM) and median (CWMed) [44], geometric median (GM) [10], and multi-Krum (MK) [7]. Useful properties of a resulting Robust-DGD method, which is summarized in Algorithm 1, are recalled below.
Robust distributed gradient descent (Robust-DGD). In Robust-DGD, robust aggregation protects the learning procedure from incorrect gradients sent by the adversarial workers. The robustness of an aggregation can be quantified by the following property of $(f,\kappa)$-robustness, proposed in [3].
@@ 3 Relevant Background
  Initialization: Server chooses a model ${{\theta}_{1}}\in\mathbb{R}^{d}$, a learning rate $\gamma\in\mathbb{R}^{+}$ and a robust aggregation rule $\mathbf{F}:\mathbb{R}^{n\times d}\to\mathbb{R}^{d}$.
@@
Aggregation rule ${\mathbf{F}:\mathbb{R}^{n\times d}\to\mathbb{R}^{d}}$ is said to be $(f,\kappa)$-robust if there exists $\kappa\in\mathbb{R}$ such that, for all $x_{1},\ldots,x_{n}\in\mathbb{R}^{d}$ and any set $S\subseteq[n]$, $\left\lvert{S}\right\rvert=n-f$, the following holds true:
@@
We refer to $\kappa$ as the robustness coefficient.
The property of $(f,\kappa)$-robustness encompasses several robust aggregations, including the ones mentioned above and more (e.g., see [3]).222$(f,\kappa)$-robustness unifies other robustness definitions: $(\delta_{\text{max}},c)$-robustness [26] and $(f,\lambda)$-resilience [14]. It has been shown that an aggregation rule is $(f,\kappa)$-robust only if $\kappa\geq\frac{f}{n-2f}$ [3]. Importantly, the asymptotic error of Robust-DGD with $(f,\kappa)$-robust aggregation, where $\kappa\in\mathcal{O}(f/(n-2f))$, matches the lower bound (recalled in Lemma 3.1[4]. Note that all the aforementioned aggregation rules, i.e., CWTM, CWMed, GM and MK, attain optimal robustness, i.e., are $(f,\mathcal{O}\left(f/(n-2f)\right))$-robust, when composed with the pre-aggregation step: nearest neighbor mixing (NNM) [3]. [^3] CWTM has been shown to be optimal even without NNM [3]. Specifically, we recall the following result.
Lemma 3.3 (Lemma 1 in [3]).
@@
Lemma 3.4 (Theorem 2 in [4]).
Consider Algorithm 1 with $\gamma\leq 1/L$. Let $\Delta_{o}\in\mathbb{R}^{+}$ such that $\mathcal{L}_{\mathcal{H}}({{\theta}_{1}})-\mathcal{L}_{\mathcal{H}}^{*}\leq\Delta_{o}$. If the aggregation rule $\mathbf{F}$ is $(f,\kappa)$-robust with $\kappa B^{2}<1$, then

### v3

# The Vital Role of Gradient Clipping in Byzantine-Resilient Distributed Learning

@@ Abstract
Byzantine-resilient distributed machine learning seeks to achieve robust learning performance in the presence of misbehaving or adversarial workers. While state-of-the-art (SOTA) robust distributed gradient descent (Robust-DGD) methods were proven theoretically optimal, their empirical success has often relied on pre-aggregation gradient clipping. However, the currently considered static clipping strategy exhibits mixed results: improving robustness against some attacks while being ineffective or detrimental against others. We address this gap by proposing a principled adaptive clipping strategy, termed Adaptive Robust Clipping (ARC). We show that ARC consistently enhances the empirical robustness of SOTA Robust-DGD methods, while preserving the theoretical robustness guarantees. Our analysis shows that ARC provably improves the asymptotic convergence guarantee of Robust-DGD in the case when the model is well-initialized. We validate this theoretical insight through an exhaustive set of experiments on benchmark image classification tasks. We observe that the improvement induced by ARC is more pronounced in highly heterogeneous and adversarial settings.
@@ 1 Introduction
Distributed machine learning, a.k.a. federated learning, has emerged as a dominant paradigm to cope with the increasing computational cost of learning tasks, mainly due to growing model sizes and datasets (Kairouz et al., 2021). Worker machines, holding each a fraction of the training dataset, collaborate over a network to learn an optimal common model over the collection of their datasets. Workers typically collaborate with the help of a central coordinator, that we call server (McMahan et al., 2017). Besides scalability, distributed learning is also helpful in preserving data ownership and sovereignty, since the workers do not have to share their local datasets during the learning.
@@
Conventional distributed learning algorithms are known to be vulnerable to misbehaving workers that could behave unpredictably (Blanchard et al., 2017; Kairouz et al., 2021; Guerraoui et al., 2023). Misbehavior may result from software and hardware bugs, data poisoning, or malicious players controlling part of the network. In the parlance of distributed computing, misbehaving workers are referred to as Byzantine (Lamport et al., 1982). Due to the growing influence of distributed learning in critical public-domain applications such as healthcare (Nguyen et al., 2022) and finance (Long et al., 2020), the problem of robustness to misbehaving workers, a.k.a. robust distributed learning, has received significant attention (Blanchard et al., 2017; Yin et al., 2018; Farhadkhani et al., 2022; Karimireddy et al., 2022; Gorbunov et al., 2023; Allouah et al., 2023a; Farhadkhani et al., 2023).
@@
Robust distributed learning algorithms primarily rely on robust aggregation, such as coordinate-wise trimmed mean (CWTM) (Yin et al., 2018), geometric median (GM) (Chen et al., 2017) and multi-Krum (MK) (Blanchard et al., 2017). Specifically, in robust distributed gradient descent (Robust-DGD), the server aggregates the workers’ local gradients using a robust aggregation method, instead of simply averaging them. This protects the learning procedure from erroneous gradients sent by misbehaving workers. Recent work has made significant improvements over these aggregation rules by incorporating a pre-aggregation step such as bucketing (Karimireddy et al., 2022; Gorbunov et al., 2023) and nearest-neighbor mixing (NNM) (Allouah et al., 2023a), to tackle gradient dissimilarity resulting from data heterogeneity. The learning guarantee of the resulting Robust-DGD has been proven to be optimal (Allouah et al., 2023b), i.e., it cannot be improved without additional assumptions under the standard heterogeneity model of $(G,B)$-gradient dissimilarity (Karimireddy et al., 2020).
Despite its theoretical tightness, the empirical success of Robust-DGD has unknowingly relied on pre-aggregation gradient clipping (Mhamdi et al., 2021; Farhadkhani et al., 2022; Allouah et al., 2023a). Specifically, clipping the gradients of the workers prior to aggregation has been observed to sometimes enhance the algorithm’s empirical performance in the presence of adversarial workers, as evidenced in Figure 1a. Yet, this improvement lacks a concrete explanation, raising the question of whether the observed benefits of clipping are merely anecdotal. This leads to the natural inquiry: Why, and when, does pre-aggregation clipping improve robustness?
(a) Sign-flipping (SF) (Allen-Zhu et al., 2020)
(b) Label-flipping (LF) (Allen-Zhu et al., 2020)
Figure 1: Impact of pre-aggregation clipping, specifically static clipping with $C=2$ and our adaptive clipping algorithm ARC, on Robust-DGD with aggregation scheme CWTM $\circ$ NNM, under the SF and LF adversarial attacks. We consider the MNIST (Deng, 2012) dataset distributed amongst $10$ honest (non-adversarial) workers with extreme heterogeneity, and there are three adversarial workers.
In particular, using a constant clipping threshold, referred to as static clipping, has exhibited mixed results. Figure 1 shows that while static clipping effectively mitigates sign-flipping (SF) attacks, it completely fails under label-flipping (LF). This indicates an inherent fragility of static clipping. Indeed, we prove in our work that static clipping breaks the standard $(f,\kappa)$-robustness property of an aggregation method (Allouah et al., 2023a). This highlights a key shortcoming of existing empirical results that rely on static clipping (Mhamdi et al., 2021; Farhadkhani et al., 2022; Allouah et al., 2023a). To overcome the limitations of static clipping but preserve its empirical benefits at the same time, we introduce a novel adaptive clipping scheme, termed Adaptive Robust Clipping (ARC).
ARC dynamically adjusts the clipping threshold as per the gradients sent by the workers and the fraction of adversarial workers to be tolerated. We demonstrate that integrating ARC into Robust-DGD consistently improves its empirical performance (see Figures 1 and 2a), while also preserving the convergence guarantee of the original Robust-DGD algorithm. Moreover, we show that when the model initialization is good, ARC provably improves the robustness of Robust-DGD. The benefits of ARC are more pronounced as the fraction of misbehaving workers approaches the system’s breakdown point [^1] Breakdown point refers the minimum fraction of adversarial workers that can break the system, making it impossible to guarantee a bound on the learning error Allouah et al. (2023b). and when the data across the workers is highly heterogeneous. Our key results are summarized below. Critical comparisons to prior work are deferred to Section 6.
Main results & contributions. We consider a system comprising $n$ workers and a server. The goal is to tolerate up to $f$ adversarial workers.
@@
(1) Adaptive robust clipping (ARC). We propose ARC, wherein prior to aggregating the gradients, the server clips the largest $k\coloneqq\left\lfloor{2(f/n)(n-f)}\right\rfloor$ gradients using a clipping parameter given by the (Euclidean) norm of the $(k+1)$-th largest gradient. It is important to note that in contrast to existing adaptive clipping schemes (Diakonikolas et al., 2020; Abdalla & Zhivotovskiy, 2024), ARC does not require additional a priori information on honest workers’ gradients. We prove that ARC preserves the robustness guarantee of the original robust aggregation method.
@@
(2) Improved empirical robustness. We conduct experiments on MNIST (Deng, 2012), Fashion-MNIST (Xiao et al., 2017), and CIFAR-10 (Krizhevsky et al., 2014), across various data heterogeneity settings and adversarial regimes. Our results demonstrate that ARC significantly enhances the performance of state-of-the-art Robust-DGD methods, particularly in scenarios with high data heterogeneity (Figure 2a) and a large number of adversarial workers (Figure 4b).
(3) Improved learning guarantee. We demonstrate that ARC possesses an additional property, we call Bounded Aggregation Output, that is not satisfied by classical robust aggregation methods. Specifically, ARC constrains the norm of an adversarial gradient by that of an honest (non-adversarial) gradient. Leveraging this property, we show that the lower bound established under data heterogeneity in Allouah et al. (2023b) can be circumvented using ARC, provided the honest workers’ gradients are bounded at model initialization. An empirical validation of this insight is shown in Figure 2b. Such model initialization is often satisfiable in practice (Glorot & Bengio, 2010), highlighting the practical relevance of ARC. When the model is arbitrarily initialized, ARC recovers the original convergence guarantee of Robust-DGD in the worst case (see Figure 2b).
(a) Improved robustness
(b) Influence of model initialization
Figure 2: Worst-case maximal accuracies of Robust-DSGD, with and without ARC, across several types of misbehavior for distributed MNIST with $10$ honest workers and under extreme heterogeneity. On the left, we vary the number of adversarial workers. On the right, we vary the initialization conditions by scaling a well-chosen set of initial parameters (CWTM $\circ$ NNM is used, and $f=1$). More details on the experimental setup can be found in Sections 4 and 5.2, and Appendix D.
2 Problem Statement and Relevant Background
[…]
@@ 2 Problem Statement and Relevant Background
A robust distributed learning algorithm aims to output a good model despite the presence of adversarial (a.k.a., Byzantine (Lamport et al., 1982)) workers in the system (Su & Vaidya, 2016; Yin et al., 2018; Allouah et al., 2023a). Specifically, the goal is to design a distributed learning algorithm that can tolerate up to $f$ adversarial workers, of a priori unknown identity, out of $n$ workers. Adversarial workers need not follow a prescribed algorithm, and can send arbitrary information to the server. In the context of distributed gradient descent, adversarial workers can send incorrect gradients to the server (Baruch et al., 2019; Allen-Zhu et al., 2020; Xie et al., 2020; Karimireddy et al., 2022). We let $\mathcal{H}\subseteq[n]$, with $\left\lvert{\mathcal{H}}\right\rvert=n-f$, denote the set of honest workers that do not deviate from the algorithm. The objective of the server is to minimize the average loss function of honest workers given by
@@
While we can find a minimum of $\mathcal{L}_{\mathcal{H}}(\theta)$ when $\mathcal{L}_{\mathcal{H}}$ is convex, in general however the loss function is non-convex, and the optimization problem is NP-hard (Boyd & Vandenberghe, 2004). Therefore, we aim to find a stationary point of $\mathcal{L}_{\mathcal{H}}$ instead, i.e., $\theta^{*}$ such that $\left\lVert{\nabla\mathcal{L}_{\mathcal{H}}(\theta^{\ast})}\right\rVert=0$. Formally, we define robustness to adversarial workers by $(f,\epsilon)$-resilience (Allouah et al., 2023a).
@@
If a distributed learning algorithm $\mathcal{A}$ is $(f,\varepsilon)$-resilient then it can tolerate up to $f$ adversarial workers. It is standard in robust distributed learning to design robust algorithms using $f$ as a parameter (Su & Vaidya, 2016; Blanchard et al., 2017; Yin et al., 2018; Karimireddy et al., 2022; Gorbunov et al., 2023; Allouah et al., 2023a), $f$ being the maximum tolerable number of adversarial workers.
In the context of distributed learning, data heterogeneity is characterized by the following standard notion of $(G,B)$-gradient dissimilarity (Vaswani et al., 2019; Karimireddy et al., 2020; 2022; Gorbunov et al., 2023; Allouah et al., 2023b).
@@
General limitations on robustness. Note that $(f,\varepsilon)$-resilience is impossible (for any $\varepsilon$) when $f/n\geq 1/2$ (Gupta & Vaidya, 2020). Moreover, even for $f/n<1/2$, it is generally impossible to achieve $(f,\varepsilon)$-resilience for arbitrarily small $\varepsilon$ due to the disparity amongst the local datasets $\mathcal{D}_{i},\;i\in\mathcal{H}$ (a.k.a., data heterogeneity) (Liu et al., 2021). Henceforth, we assume that $n>2f$. Under $(G,B)$-gradient dissimilarity, we have the following lower bound on the achievable resilience.
Lemma 2.3 (Non-convex extension of Theorem 1 Allouah et al.).
Under $(G,B)$-gradient dissimilarity, a distributed learning algorithm is $(f,\varepsilon)$-resilient only if $\frac{f}{n}<\frac{1}{2+B^{2}}$ and $\varepsilon\geq\frac{1}{4}\cdot\frac{f}{n-(2+B^{2})f}G^{2}$.
For a given distributed learning problem, the minimum fraction of adversarial workers that renders any distributed learning algorithm ineffective is called the breakdown point (Guerraoui et al., 2023; Allouah et al., 2023b). Thus, according to Lemma 2.3, there exists a distributed learning problem satisfying $(G,B)$-gradient dissimilarity whose breakdown point is given by $1/(2+B^{2})$.
Robust distributed gradient descent. To tolerate adversarial workers, we replace the averaging operation in the classic distributed gradient descent (DGD) method with a robust aggregation, e.g., coordinate-wise trimmed mean (CWTM) and median (CWMed) (Yin et al., 2018), geometric median (GM) (Chen et al., 2017), and multi-Krum (MK) (Blanchard et al., 2017). This yields Robust-DGD, presented in Algorithm 1, where robust aggregation protects the learning from incorrect gradients sent by the adversarial workers. The robustness of an aggregation method can be quantified by the following property of $(f,\kappa)$-robustness, proposed by Allouah et al. (2023a).
@@
  Initialization: Server chooses a model ${{\theta}_{1}}\in\mathbb{R}^{d}$, a learning rate $\gamma\in\mathbb{R}^{+}$ and a robust aggregation method $\mathbf{F}:\mathbb{R}^{n\times d}\to\mathbb{R}^{d}$.
@@
${\mathbf{F}:\mathbb{R}^{n\times d}\to\mathbb{R}^{d}}$ is said to be $(f,\kappa)$-robust if there exists a robustness coefficient $\kappa\in\mathbb{R}$ such that for all $x_{1},\ldots,x_{n}\in\mathbb{R}^{d}$ and any set $S\subseteq[n]$, $\left\lvert{S}\right\rvert=n-f$, the following holds:
@@
$(f,\kappa)$-robustness [^2] Unifies $(\delta_{\text{max}},c)$-robustness (Karimireddy et al., 2021) and $(f,\lambda)$-resilience (Farhadkhani et al., 2022). encompasses several robust aggregations, including the ones mentioned above and more (e.g., see Allouah et al. (2023a)). It has been shown that an aggregation method is $(f,\kappa)$-robust only if $\kappa\geq\frac{f}{n-2f}$ (Allouah et al., 2023a). Importantly, the asymptotic error of Robust-DGD with $(f,\kappa)$-robust aggregation, where $\kappa\in\mathcal{O}(f/(n-2f))$, matches the lower bound (recalled in Lemma 2.3(Allouah et al., 2023b). Note that the aforementioned aggregation methods (CWTM, CWMed, GM and MK) attain optimal robustness when composed with the pre-aggregation scheme: nearest neighbor mixing (NNM) (Allouah et al., 2023a). [^3] CWTM has been shown to be optimal even without NNM. Specifically, we recall the following result.
Lemma 2.5 (Lemma 1 in Allouah et al. (2023a)).
@@
Lemma 2.6 (Theorem 2 in Allouah et al. (2023b)).
Consider Algorithm 1 with $\gamma\leq 1/L$. Let $\Delta_{o}\in\mathbb{R}^{+}$ such that $\mathcal{L}_{\mathcal{H}}({{\theta}_{1}})-\mathcal{L}_{\mathcal{H}}^{*}\leq\Delta_{o}$. If $\mathbf{F}$ is $(f,\kappa)$-robust with $\kappa B^{2}<1$, then